# On Vanishing Gradients, Over-Smoothing, and Over-Squashing in GNNs: Bridging Recurrent and Graph Learning

### v1

@@ 5 The Impact of Vanishing Gradients on Over-squashing
In this section, we study the connection between vanishing gradients and over-squashing in GNNs.
@@ 5.1 Mitigating over-squashing by combining increased connectivity and non-dissipativity
Theorem 5.1 (Sensitivity bounds, Di Giovanni et al. (2023)).
@@
with $\mathbf{O}=c_{r}\mathbf{I}+c_{a}\mathbf{A}\in\mathbb{R}^{n\times n}$ is the message passing matrix adopted by the MPNN, with $c_{r}$ and $c_{a}$ are the contributions of the self-connection and aggregation term.
Theorem 5.1 shows that the sensitivity of the node embedding arises from a combination of (i) a term based on the graph topology and (ii) a term dependent on the model dynamics, with over-squashing occurring when the right-hand side of Equation (15) becomes too small. We highlight that this differs from the standard product Jacobian which arises in RNNs. This is because in MPNNs, messages are scaled by the inverse node degree, incurring an extra information dissipation step. Consequently, while recurrent architectures only need to adjust their dynamics to ensure long memory, MPNNs must simultaneously enhance graph connectivity and modify their dynamics to mitigate vanishing gradients.
Even though the sensitivity bound in Theorem 5.1 is controlled by two components, the majority of the literature has typically focused on addressing only the topological term via graph rewiring (Gasteiger et al., 2019; Topping et al., 2021; Karhadkar et al., 2022; Barbero et al., 2023; Finkelshtein et al., 2024), with some methods also targeting the model dynamics (Gravina et al., 2023, 2025; Heilig et al., 2025). In fact, Di Giovanni et al. (2023) explicitly discourages increasing the model term in Theorem 5.1 and claims that doing so could lead to over-fitting and poorer generalization. However, we argue that increasing the model term — directly linked to vanishing gradients as discussed in Section 4 — is essential to mitigate over-squashing. Rather than harming performance, boosting this term helps prevent over-smoothing, since even in a well-connected graph where information can be reached in fewer hops, unaddressed vanishing gradients due to the model term will cause the target node’s features to decay to zero during message passing.
Frameworks combining these strategies include Gutteridge et al. (2023), which integrates graph rewiring with a delay term, and Ding et al. (2024), which merges multi-hop aggregation with ideas from SSMs. [^6] Further links between the delay term and vanishing gradients are discussed in Appendix D.3. Further, we show that models tend to converge to the edge of chaos during training in Appendix D.4. These approaches have generally led to state-of-the-art results, significantly improving performance over standalone rewiring techniques.
@@ 5.2 Empirical validation of claims
Figure 5: Left: Performance on the RingTransfer task for different models. Right: Effect of dissipativity on performance.
Table 1: Ablation on LRGB datasets. Here, $d\uparrow$ means and increase in the latent dimention, while -- indicates the removal of a component and $+$ indicates an addition of a component.
Model Pept-func Pept-struct
AP $\uparrow$ MAE $\downarrow$
GCN $60.93_{\pm 0.138}$ $33.41_{\pm 0.041}$
kGCN-SSM $\underline{69.02}_{\pm 0.218}$ $28.98_{\pm 0.324}$
+ $d\uparrow$ $\mathbf{72.12}_{\pm 0.268}$ $27.01_{\pm 0.071}$
-- $\text{eig}(\Lambda)\approx 1$ $61.41_{\pm 0.724}$ $\mathbf{25.81}_{\pm 0.032}$
-- SSM $57.76_{\pm 1.971}$ $\underline{26.02}_{\pm 0.213}$
-- khop $60.93_{\pm 0.138}$ $33.41_{\pm 0.041}$
DRew-GCN $68.04_{\pm 1.442}$ $27.66_{\pm 0.187}$
+ $d\uparrow$ $68.05_{\pm 0.626}$ $27.64_{\pm 0.067}$
-- Delay $49.02_{\pm 2.512}$ $27.08_{\pm 0.041}$
We start by testing the performance on the RingTransfer task introduced in Di Giovanni et al. (2023), as it is a task where we certifiably know that long-range dependencies exist. We modify the $\text{eig}(\Lambda)$ in the kGNN-SSM to move the Jacobian from the edge of stability to a progressively more dissipative state. The results are shown in Figure 5. From the figure, we see that (i) kGNN-SSM achieves state-of-the art performance only when coupling strong connectivity and an edge of chaos Jacobian (ii) making the model more dissipative directly results in worse long-range modeling capabilities. We believe the latter point demonstrates the importance of the model term in Theorem 5.1.
Next, we ablate each component of the model on three graph property prediction tasks introduced in Gravina et al. (2023) alongside the real-world long-range graph benchmark (LRGB) from Dwivedi et al. (2022), focusing on the peptides-func and peptides-struct tasks. Additional details regarding the datasets and the experimental setting are reported in Appendix C. Here, we focus on ablating the effect of rewiring, adding an SSM layer, and placing the model at the edge of chaos through $\mathbf{\Lambda}$. In the LRBG tasks, we additionally ablate the effect of increasing the hidden memory size, as we consider forty layers in the peptides-func dataset, which requires more long-range capabilities. Here, we also ablate DRew (Gutteridge et al., 2023) under the same settings. We also provide a more detailed comparison with other models in Appendix D.4, and provide additional comments around the LRGB tasks in Appendix D.5.
Table 2: Mean and std. for test $log_{10}(\mathrm{MSE})$ averaged over 4 random weight initializations on the GPP tasks. Lower is better.
@@
+ $\text{eig}(\Lambda)\approx 1$2.444±0.098 -3.593±0.103 -2.258±0.009
@@
The results are shown in Tables 1 and 2. Across the board, we observe that kGNN-SSM not only matches DRew-Delay, but also outperforms it by a large amount in all cases, showcasing the strength of our state-space approach. In particular, we generally observe significant decreases in performance when removing both the high connectivity and non-dissipativity components of the model, highlighting their individual importance. Finally, we see that increasing memory size plays a big role in the peptides-struct task, which is in line with what has been observed in sequence modeling (Gu et al., 2021).
Message of the Section: Oversquashing in GNNs arises from both graph connectivity and the model’s capacity to avoid vanishing gradients. While most studies focus on connectivity, we argue that preserving signal strength through non-dissipative model dynamics is equally important. High connectivity allows nodes of interest to be reached in fewer message-passing steps while model dynamics ensure information is preserved.
@@ 6 Conclusion
In this work, we revisit the well-known problems of over-smoothing and over-squashing in GNNs from the lens of vanishing gradients, by studying GNNs from the perspective of recurrent and state-space models. In particular, we show that GNNs are prone to a phenomenon we term extreme gradient vanishing, which results in ill-conditioned signal propagation with few layers. As such, we argue that it is important to control the layerwise Jacobian and propose a state-space-inspired GNN model, termed GNN-SSM, to do so. We then uncover that vanishing gradients result in a very specific form of over-smoothing in which all signals converge exactly to the $\mathbf{0}$ vector, and support this claim empirically. Finally, we theoretically argue and empirically show that mitigation of over-squashing is best achieved through a combination of strong graph connectivity and non-dissipative model dynamics.
Limitations and Future Work.
We believe our work opens up a number of interesting directions that aim to bridge the gap between graph and sequence modeling. In particular, we hope that this work will encourage researchers to adapt vanishing gradient mitigation methods from the sequence modeling community to GNNs, and conversely explore how graph learning ideas can be brought to recurrent models. In our work, we mostly focused on GCN and GAT type updates, but we believe that our analysis can be extended to understand how different choices of updates and non-linearities affect training dynamics, which we leave for a future work.
Acknowledgements
AA and XD thank the Oxford-Man Institute for financial support. AA thanks T. Anderson Keller for engaging comments on early versions of the manuscript, Max Welling for valuable pointers to the dynamical isometry literature, and T. Konstantin Rusch for sharing the code used for the Dirichlet energy experiments.
Impact Statement
This work advances the study of Graph Neural Networks (GNNs) by establishing a theoretical link between vanishing gradients, over-smoothing, and over-squashing in Message Passing Neural Networks (MPNNs). We propose a principled framework that investigates whether a common underlying cause governs these phenomena, alongside a state-space-inspired GNN model that effectively mitigates them. The research presented herein contributes to the broader understanding and development of GNNs. In this work, we do not release any datasets or models that could pose a significant risk of misuse. We believe our research does not have any direct or indirect negative societal implications or harmful consequences. As far as we are aware, this study does not raise any ethical concerns or potential negative impacts.
@@ A.1 Proofs of Jacobian Theorems
Definition A.3 (Marchenko–Pastur distribution. (Marchenko & Pastur, 1967)).
In the high-dimensional limit ($n,p\to\infty$ at a fixed ratio $p/n\to c$), the empirical eigenvalue distribution of the (properly normalized) Wishart matrix $\mathbf{X}^{T}\mathbf{X}$ converges to the Marchenko–Pastur distribution. Concretely, if $\mathbf{X}\in\mathbb{R}^{n\times p}$ has entries $\mathcal{N}(0,1)$, then the eigenvalues of $\mathbf{X}^{T}\mathbf{X}$ lie within $[(1-\sqrt{c})^{2},\,(1+\sqrt{c})^{2}]$ for large $n,p$, and their density converges to
@@ A.2 Proofs to Smoothing Theorems
The notion of Lipschitz continuity is effectively a bound on the rate of change of a function. It is therefore not surprising that one can relate the Lipschitz constant to the Jacobian of $f$. In particular, we state a useful and well-known result (Hassan et al., 2002) that relates the (continuous) Jacobian map $\mathbf{J}_{f}$ of a continuous function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ to its Lipschitz constant $L\geq 0$. In particular, the Lipschitz constant is is the supremum of the (induced) norm of the Jacobian taken over its domain.
Lemma A.8 (Hassan et al. (2002)).
@@
The condition of $U$ being convex is a technicality that is easily achieved in practice with the assumption that input features are bounded and that therefore they live in a convex hull $U$. In particular, at each layer $k$ one can also find a convex hull $U_{k}$ such that the image of the layer $k-1$ is contained within $U_{k}$. We highlight that for non-linearities such as ReLU, there are technical difficulties when taking this supremum as there is a non-differentiable point at 00. This can be circumvented by considering instead a supremum of the (Clarke) generalized Jacobian (Jordan & Dimakis, 2020). We ignore this small detail in this work for simplicity as for ReLU this is equivalent to considering the supremum over $U/\mathbf{0}$, i.e. simply ignoring the problematic point $\mathbf{0}$.
@@
It is then clear that, if $\left\lVert{f_{\ell}}\right\rVert_{\text{Lip}}\leq 1-\epsilon$ for some $\epsilon>0$ for all $\ell=1\dots L$, $\prod_{\ell=1}^{L}\left\lVert{f_{\ell}}\right\rVert_{\text{Lip}}^{2}\leq(1-\epsilon)^{2L}\to 0$ as $L\to\infty$. ∎
@@
and $d_{G}:V\times V\rightarrow\mathbb{R}_{\geq 0}$ is the length of the minimal walk connecting nodes $i$ and $j$. This approach avoids a large amount of information being squashed into a single vector, and is more in line with the recurrent paradigm. We note that this scheme is similar to (Ding et al., 2024), but in this case we do not consider different block or parameter sharing, and our recurrent mechanism is based on an untrained SSM layer.
@@ Appendix C Experimental Details
In this task, we aim to analyze the dynamics of the Dirichlet energy across three different graph topologies: Cora (Yang et al., 2016), Texas (Pei et al., 2020), and a grid graph. The Cora dataset is a citation network consisting of 2,708 nodes (papers) and 10,556 edges (citations). The Texas dataset represents a webpage graph with 183 nodes (web pages) and 499 edges (hyperlinks). Lastly, the grid graph is a two-dimensional $10\times 10$ regular grid with 4-neighbor connectivity. For all three graphs, node features are randomly initialized from a normal distribution with a mean of 0 and variance of 1. These node features are then propagated over 80 layers (or iterations) using untrained GNNs to observe the energy dynamics.
@@
We employ the same experimental setting and data outlined in (Gravina et al., 2023). Each model is designed as three components: the encoder, the graph convolution, and the readout. We perform hyperparameter tuning via grid search, optimizing the Mean Square Error (MSE). The models are trained using the Adam optimizer for a maximum of 1500 epochs, with early stopping based on the validation error, applying a 100 epochs patience. For each model configuration, we perform 4 training runs with different weight initializations and report the average results. We report in Table 3 the employed grid of hyperparameters.
@@
We consider the peptides-func and peptides-struct datasets from (Dwivedi et al., 2022). Both datasets consist of 15,535 graphs, where each graph corresponds to 1D amino acid chain (i.e., peptide), where nodes are the heavy atoms of the peptide and edges are the bonds between them. peptides-func is a multi-label graph classification dataset whose objective is to predict the peptide function, such as antibacterial and antiviral function. peptides-struct is a multi-label graph regression dataset focused on predicting the 3D structural properties of peptides, such as the inertia of the molecule and maximum atom-pair distance.
We use the same experimental setting and splits from (Dwivedi et al., 2022). We perform hyperparameter tuning via grid search, optimizing the Average Precision (AP) in the Peptide-func and Mean Absolute Error (MAE) in the Peptide-struct. The models are trained using the AdamW optimizer for a maximum of 300 epochs. For each model configuration, we perform four training runs with different weight initializations and report the average results. We report in Table 3 the employed grid of hyperparameters.

### v2

@@ 5 The Impact of Vanishing Gradients on Over-squashing
In this section, we study the connection between vanishing gradients and over-squashing in GNNs, which fills in the gaps and open questions left in the analysis of (di2023over, ).  [^9] In this part, we mostly focus on long-range interactions, which has been in many cases treated synonymously with over-squashing in the literature. For a more precise characterization of the relationship between over-squashing and long range, we point the reader to arnaiz2025oversmoothing . We highlight that we believe the distinction between computational and topological bottlenecks is not particularly relevant in our setting.
@@ 5.1 Mitigating over-squashing by combining increased connectivity and non-dissipativity
Theorem 5.1 (Sensitivity bounds, (di2023over, )).
@@
where $\mathbf{O}=c_{r}\mathbf{I}+c_{a}\mathbf{A}\in\mathbb{R}^{n\times n}$ is the message passing matrix adopted by the MPNN, and where $c_{r}$ and $c_{a}$ are the contributions of the self-connection and aggregation term.
Theorem 5.1 shows that the sensitivity of the node embedding is a combination of (i) a term based on the graph topology and (ii) a term dependent on the model dynamics, with over-squashing occurring when the right-hand side of Equation (14) becomes too small. We highlight that this differs from the standard product Jacobian, which arises in RNNs. This is because in MPNNs, messages are scaled by the inverse node degree, incurring an extra information dissipation step. Consequently, while recurrent architectures only need to adjust their dynamics to ensure long memory, MPNNs must simultaneously enhance graph connectivity and modify their dynamics to mitigate vanishing and exploding gradients.
Table 1: Ablation on LRGB datasets. $d\uparrow$ adds latent dims; - removes; ++ adds.
Model Pept-func Pept-struct
[…]
GCN $60.93_{\pm 0.138}$ $33.41_{\pm 0.041}$
kGCN-SSM $\underline{69.02}_{\pm 0.218}$ $28.98_{\pm 0.324}$
$d\uparrow$ $\mathbf{72.12}_{\pm 0.268}$ $27.01_{\pm 0.071}$
$\mathrm{eig}(\Lambda)\approx 1$ $61.41_{\pm 0.724}$ $\mathbf{25.81}_{\pm 0.032}$
$\mathrm{SSM}$ $57.76_{\pm 1.971}$ $\underline{26.02}_{\pm 0.213}$
$\mathrm{khop}$ $60.93_{\pm 0.138}$ $33.41_{\pm 0.041}$
DRew-GCN $68.04_{\pm 1.442}$ $27.66_{\pm 0.187}$
$d\uparrow$ $68.05_{\pm 0.626}$ $27.64_{\pm 0.067}$
$\mathrm{Delay}$ $49.02_{\pm 2.512}$ $27.08_{\pm 0.041}$
Even though the sensitivity bound in Theorem 5.1 is controlled by two components, the majority of the literature has typically focused on addressing only the topological term via graph rewiring diffusion_improves ; topping2021understanding ; karhadkar2022fosr ; barbero2023locality ; finkelshteincooperative2024 , with some methods also targeting the model dynamics gravina2022anti ; gravina_swan ; heilig2024injecting . In fact, (di2023over, ) explicitly discourages increasing the model term in Theorem 5.1 and claims that doing so could lead to over-fitting and poorer generalization. However, we argue that increasing the model term, directly linked to vanishing gradients as discussed in Section 4, is essential to mitigate over-squashing. Rather than harming performance, boosting this term helps prevent over-smoothing, since even in a well-connected graph where information can be reached in fewer hops, unaddressed vanishing gradients due to the model term will cause the target node’s features to collapse during message passing. Frameworks combining these strategies include (gutteridge2023drew, ), which integrates graph rewiring with a delay term, and (ding2024recurrent, ), which merges multi-hop aggregation with ideas from SSMs. [^10] Further links between the delay term and vanishing gradients are discussed in Appendix D.2. Further, we show that models tend to converge to the edge of chaos during training in Appendix D.3. These approaches have generally led to state-of-the-art results, significantly improving performance over standalone rewiring.
@@ 5.2 Empirical validation of claims
Figure 5: Left: Performance on the RingTransfer task. Right: Effect of dissipativity.
We start by testing the performance on the RingTransfer task introduced in (di2023over, ), as it is a task where we certifiably know that long-range dependencies exist. We modify the $\text{eig}(\Lambda)$ in the kGNN-SSM to move the Jacobian from the edge of stability to a progressively more dissipative state. The results are shown in Figure 5. From the figure, we see that (i) kGNN-SSM achieves state-of-the art performance only when coupling strong connectivity and an edge of chaos Jacobian (ii) making the model more dissipative directly results in worse long-range modeling capabilities. We believe the latter point demonstrates the importance of the model term in Theorem˜5.1.
Table 2: Mean and std. for test $log_{10}(\mathrm{MSE})$ averaged over 4 random weight initializations on the GPP tasks
@@
+ $\text{eig}(\Lambda)\approx 1$ -2.444±0.098 -3.593±0.103 -2.258±0.009
@@
Next, we ablate each component of the model on three graph property prediction tasks introduced in (gravina2022anti, ) alongside the real-world long-range graph benchmark (LRGB) from (dwivedi2022LRGB, ), focusing on the peptides-func and peptides-struct tasks. Additional details regarding the datasets and the experimental setting are reported in Appendix C. Here, we focus on ablating the effect of rewiring, adding an SSM layer, and placing the model at the edge of chaos through $\mathbf{\Lambda}$. In the LRBG tasks, we additionally ablate the effect of increasing the hidden memory size, as we consider forty layers in the peptides-func dataset, which requires more long-range capabilities. Here, we also ablate DRew gutteridge2023drew under the same settings. We also provide a more detailed comparison with other models in Appendix D.3, and provide additional comments around the LRGB tasks in Appendix D.4.
The results are shown in Tables 1 and 2. Across the board, we observe that kGNN-SSM not only matches DRew-Delay, but also outperforms it by a large amount in all cases, showcasing the strength of our state-space approach. In particular, we generally observe significant decreases in performance when removing both the high connectivity and non-dissipativity components of the model, highlighting their individual importance. Finally, we see that increasing memory size plays a big role in the peptides-func task, which is in line with observations made the in sequence modeling literature gu2021 .
Message of the Section: The inability of GNN models to retrieve information from distant nodes arises from both graph connectivity and the model’s capability to avoid vanishing and exploding gradients. While most studies focus on connectivity, we argue that preserving signal strength through non-dissipative model dynamics is equally important. High connectivity allows nodes of interest to be reached in fewer message-passing steps while model dynamics ensure information is preserved.
@@ 6 Conclusion
In this work, we revisit the well-known problems of representational collapse and over-squashing in GNNs from the lens of vanishing gradients, by studying GNNs from the perspective of recurrent and state-space models. In particular, we show that GNNs are prone to a phenomenon we term extreme gradient vanishing, which results in ill-conditioned signal propagation with few layers. As such, we argue that it is important to control the layerwise Jacobian and propose a state-space-inspired GNN model, termed GNN-SSM, to do so. We then uncover that vanishing gradients result in a specific form of over-smoothing in which all signals converge exactly to a unique fixed point, and support this claim empirically. Finally, we theoretically argue and empirically show that the mitigation of over-squashing is best achieved through a combination of strong graph connectivity and non-dissipative dynamics. Impact Statement. We aim to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments and Disclosure of Funding
AA and XD thank the Oxford-Man Institute for financial support. XD also acknowledges support from EPSRC No. EP/T023333/1. AA thanks T. Anderson Keller for engaging comments on early versions of the manuscript, Max Welling for valuable pointers to the dynamical isometry literature. AG and CG acknowledge funding from EU-EIC EMERGE (Grant No. 101070918). CG acknowledges support from NEURONE, a project funded by the European Union - Next Generation EU, M4C1 CUP I53D23003600006, under program PRIN 2022 (prj. code 20229JRTZA).
@@ A.1 Proofs of Jacobian Theorems
Definition A.3 (Marchenko–Pastur distribution. [72]).
In the high-dimensional limit ($n,p\to\infty$ at a fixed ratio $p/n\to c$), the empirical eigenvalue distribution of the (properly normalized) Wishart matrix $\mathbf{X}^{T}\mathbf{X}$ converges to the Marchenko–Pastur distribution. Concretely, if $\mathbf{X}\in\mathbb{R}^{n\times p}$ has entries $\mathcal{N}(0,1)$, then the eigenvalues of $\mathbf{X}^{T}\mathbf{X}$ lie within $(1-\sqrt{c})^{2},\,(1+\sqrt{c})^{2}]$ for large $n,p$, and their density converges to
@@ A.2 Proofs to Smoothing Theorems
The notion of Lipschitz continuity is effectively a bound on the rate of change of a function. It is therefore not surprising that one can relate the Lipschitz constant to the Jacobian of $f$. In particular, we state a useful and well-known result [58] that relates the (continuous) Jacobian map $\mathbf{J}_{f}$ of a continuous function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ to its Lipschitz constant $L\geq 0$. In particular, the Lipschitz constant is is the supremum of the (induced) norm of the Jacobian taken over its domain.
Lemma A.8 ([58]).
@@
The condition of $U$ being convex is a technicality that is easily achieved in practice with the assumption that input features are bounded and that therefore they live in a convex hull $U$. In particular, at each layer $k$ one can also find a convex hull $U_{k}$ such that the image of the layer k1k-1 is contained within $U_{k}$. We highlight that for non-linearities such as ReLU, there are technical difficulties when taking this supremum as there is a non-differentiable point at 0. This can be circumvented by considering instead a supremum of the (Clarke) generalized Jacobian [64]. We ignore this small detail in this work for simplicity as for ReLU this is equivalent to considering the supremum over $U/\mathbf{0}$, i.e. simply ignoring the problematic point $\mathbf{0}$.
@@
It is then clear that, if $\left\lVert{f_{\ell}}\right\rVert_{\text{Lip}}\leq 1-\epsilon$ for some $\epsilon>0$ for all $\ell=1\dots L$, $\prod_{\ell=1}^{L}\left\lVert{f_{\ell}}\right\rVert_{\text{Lip}}^{2}\leq(1-\epsilon)^{2L}\to 0$ as $L\to\infty$.
We note that a similar procedure was used in [80, 18] for the specific case of GCNs. Our procedure is more general, as we use the Lipschitz constant of the network, which only requires knowledge of the input-output Jacobian of each layer of the network. In the case of GCN, this would encapsulate the dynamics of the adjacency and weight matrix, and also allows us to understand how any GNN (no matter how complex its internal structure) affects the Dirichlet energy, without requiring the use of heavy assumptions or simplifications for mathematical tractability. ∎
@@
and $d_{G}:V\times V\rightarrow\mathbb{R}_{\geq 0}$ is the length of the minimal walk connecting nodes $i$ and $j$. This approach avoids a large amount of information being squashed into a single vector, and is more in line with the recurrent paradigm. We note that this scheme is similar to [32], but in this case we do not consider different block or parameter sharing, and our recurrent mechanism is based on an untrained SSM layer.
@@ Appendix C Experimental Details
In this task, we aim to analyze the dynamics of the Dirichlet energy across three different graph topologies: Cora [108], Texas [83], and a grid graph. The Cora dataset is a citation network consisting of 2,708 nodes (papers) and 10,556 edges (citations). The Texas dataset represents a webpage graph with 183 nodes (web pages) and 499 edges (hyperlinks). Lastly, the grid graph is a two-dimensional $10\times 10$ regular grid with 4-neighbor connectivity. For all three graphs, node features are randomly initialized from a normal distribution with a mean of 0 and variance of 1. These node features are then propagated over 80 layers (or iterations) using untrained GNNs to observe the energy dynamics.
@@
We employ the same experimental setting and data outlined in [47]. Each model is designed as three components: the encoder, the graph convolution, and the readout. We perform hyperparameter tuning via grid search, optimizing the Mean Square Error (MSE). The models are trained using the Adam optimizer for a maximum of 1500 epochs, with early stopping based on the validation error, applying a 100 epochs patience. For each model configuration, we perform 4 training runs with different weight initializations and report the average results. We report in Table 3 the employed grid of hyperparameters.
@@
We consider the peptides-func and peptides-struct datasets from [36]. Both datasets consist of 15,535 graphs, where each graph corresponds to 1D amino acid chain (i.e., peptide), where nodes are the heavy atoms of the peptide and edges are the bonds between them. peptides-func is a multi-label graph classification dataset whose objective is to predict the peptide function, such as antibacterial and antiviral function. peptides-struct is a multi-label graph regression dataset focused on predicting the 3D structural properties of peptides, such as the inertia of the molecule and maximum atom-pair distance.
We use the same experimental setting and splits from [36]. We perform hyperparameter tuning via grid search, optimizing the Average Precision (AP) in the Peptides-func and Mean Absolute Error (MAE) in the Peptide-struct. The models are trained using the AdamW optimizer for a maximum of 300 epochs. For each model configuration, we perform four training runs with different weight initializations and report the average results. We report in Table 3 the employed grid of hyperparameters.